OpenStudy (anonymous):

SALE PRICE: a department store is having a storewide twenty percent discount sale. the sale price S of an item that has a regular price of R is S=R-0.2R write the inverse equation. what is the regular price for an item that is on sale for $38.40??????????

OpenStudy (kropot72):

S = R - 0.2R Can you simplify the right hand side?

OpenStudy (anonymous):

I'm confuzzled :/

OpenStudy (kropot72):

What do you get when you subtract 0.2R from R? R - 0.2R = ?

OpenStudy (kropot72):

@jerr1996 Are you there?

OpenStudy (anonymous):

sorry computer problems :/ wouldn't that leave you .2?

OpenStudy (kropot72):

No. Just to review the principle involved what do you get when you subtract R from 10R? 10R - R = ?

OpenStudy (anonymous):

9R?

OpenStudy (kropot72):

Correct! So can you try this one again: R - 0.2R = ?

OpenStudy (anonymous):

.8r?

OpenStudy (kropot72):

Good work! So now the original equation is changed as follows: S = R - 0.2R = 0.8R S = 0.8R ...............(1) Now we need to find the regular price R in terms of the sale price S. To do this you divide both sides of equation (1) by 0.8. Can you try that?

OpenStudy (anonymous):

would dividing it make it a fraction?

OpenStudy (kropot72):

Yes, you get a fraction on the left hand side but not on the right hand side.

OpenStudy (anonymous):

so would it be S/.8=R?

OpenStudy (kropot72):

Good work again :) So now you have arrived at the 'inverse equation' and can calculate the regular price R by dividing the sale price (S = $38.40) by 0.8.

OpenStudy (anonymous):

48?

OpenStudy (kropot72):

R = $48.00 is correct.

OpenStudy (anonymous):

Ahhhh!!!! :D Thank You!!! :D

OpenStudy (kropot72):

You're welcome :)